OpenStudy (anonymous):

9 to the 3 power times 9 = 9 to the 7 power

OpenStudy (destinymasha):

9^3 x 9 = 9^7 ?

OpenStudy (destinymasha):

9^3 x 9 = 6561, 9^7 = 4782969 so thats false

OpenStudy (destinymasha):

Thanks for the medal!